OpenStudy (anonymous):

To solve 0.4x-0.2=1 what are the steps some one help me

OpenStudy (anonymous):

you have to add -0.2 to both sides so it looks like this 0.4x-0.2=1 +0.2 +0.2 0.4x=1+0.2 add 1 and 0.2 = 1.2 0.4x=1.2 divide 0.4 from both sides and you'll get 0.3333

OpenStudy (anonymous):

Thanks omg youre very helpful

OpenStudy (anonymous):

no problem